To a 25 ml H 2 O 2 solution, excess of acidified solution of potassium iodide was added. The iodine liberated required 20 ml of 0.3 N sodium thiosulphate solution. Calculate the volume strength of H 2 O 2 solution.

To a 25 ml H2O2 solution, excess of acidified solution of potassium iodide was added. The iodine liberated required 20 ml of 0.3 N sodium thiosulphate solution. Calculate the volume strength of H2O2 solution.

Grade:10

1 Answers

Jitender Pal
9 years ago
Meq. Of H2O2 = Meq. Of Na2S2O3
w/17 * 1000 = 20 * 0.3 ∴ w = 0.102 g
H2O2 → H2O + 1/2 O2
∴ Volume of O2 = 11200 *0.102/34 = 33.6 mL
∴ Volume strength = 33.6/25 = 1.344